harmonic conjugate of u(x,y)=e^y cosx is
1
Expert's answer
2021-07-26T16:34:35-0400

By using one of the Cauchy Riemann equations we have "\frac{\partial v}{\partial x}=-\frac{\partial u}{\partial y}=-e^y\cos x," and partially integrating with respect to "x" gives "v=-e^y\sin x+C(y)." Partially differentiating this expression and using the otherCauchy Riemann equation gives "\frac{\partial v}{\partial y}=-e^y\sin x+C'(x)=\frac{\partial u}{\partial x}=-e^y\sin x." It follows that "C'(x)=0," and hence "C(x)=C" is constant. We conclude that harmonic conjugate of "u(x,y)=e^y \cos x" is "v=-e^y\sin x+C."